OpenStudy (anonymous):

Help please and thank you. (This is pretty long and I would need help with the steps I am missing) ty!

OpenStudy (anonymous):

Simplify. \[\frac{ x+1 }{ x^2+x-12 }-\frac{ x-3 }{ x^2+7x+12 }\]

OpenStudy (anonymous):

x^2+x-12 is (x-3) (x+4) and x^2+7x+12 is (x+3)(x+4)

OpenStudy (anonymous):

so far so good

OpenStudy (anonymous):

After this I need help

OpenStudy (anonymous):

to clean it up, let a = (x-3) b = (x+4) c = (x+3) \[\frac{ x+1 }{ ab }-\frac{ x-3 }{bc}\] how do we add fractions?

OpenStudy (anonymous):

LCM is next but I'm lost

OpenStudy (anonymous):

I need help understanding how to get the LCM in this equation

OpenStudy (anonymous):

\[\frac cc\frac{ x+1 }{ ab }-\frac aa\frac{ x-3 }{bc}\] \[\frac{ c(x+1)-a(x-3) }{ abc }\] but we defined x-3 as a \[\frac{ c(x+1)-aa }{ abc }\] \[\frac{ c(x+1)-a^2 }{ abc }\] do you have to expand it all out?

OpenStudy (anonymous):

ah, the lcm would be the product of the denominators. abc

OpenStudy (anonymous):

Using this method of cleaning it up I am unable to grasp

OpenStudy (anonymous):

do we know the value of (x-3) ?

OpenStudy (anonymous):

My answer was 10x-21/(x+4)(x+3)(x-3) but that is incorrect

OpenStudy (anonymous):

the bottom looks good lets work on the top

OpenStudy (anonymous):

-2+12/(x+4)(x+3)(x-3) is incorrect as well that was my second attempt

OpenStudy (anonymous):

c(x+1) - a^2 (x+3)(x+1) - (x-3)^2 (x^2+x+3x+3) - (x^2-6x+9) x^2 +4x +3 -x^2 +6x - 9 ------------- 0 +10x-6

OpenStudy (anonymous):

meant to say -2x+12**/"()()()"

OpenStudy (anonymous):

yeah, the top works out to be 10x-6

OpenStudy (anonymous):

by clean up i meant that we keep the same terms, but just dont clutter the space with all that (x+..)(x-....) typing. its the objects the get moved about so what we name them is not important was all

OpenStudy (anonymous):

I understand I just don't follow that and what you wrote before where you get 0+10x-6 I am trying to follow how you go there the process

OpenStudy (anonymous):

there was also a part you had aa one is a and the other a represents x-3 I'm just completely lost

OpenStudy (anonymous):

since i defined a as (x-3) i got to a point of a(x-3), which is just a, times a ... a^2. or (x-3)^2

OpenStudy (anonymous):

spose we dont clean it up ....

OpenStudy (anonymous):

theres jsut more room for typing in an error is all ... \[\frac{ x+1 }{ x^2+x-12 }-\frac{ x-3 }{ x^2+7x+12 }\] \[\frac{ (x+1) }{(x-3)(x+4) }-\frac{ (x-3) }{(x+3)(x+4) }\] \[\frac{ (x+1)(x+3) }{(x-3)(x+3)(x+4) }-\frac{ (x-3)(x-3) }{(x-3)(x+3)(x+4) }\]